OpenStudy (anonymous):

What is the solution to the equation? 6 + 2s – 8s = 18 A. s = –4 B. s = –2 C. s = 2 D. s = 4

OpenStudy (anonymous):

@jadeybells @beccaboo333

OpenStudy (anonymous):

b

OpenStudy (anonymous):

6 + 2s - 8s = 18 -6 -6 -6s = 12 s = -2

OpenStudy (anonymous):

6-6s=18 12=-6s s=-2

OpenStudy (anonymous):

Thanks both of you!

OpenStudy (anonymous):

yep

OpenStudy (paki):

-2